# Sharp Capacity Scaling of Spectral Optimizers in Learning Associative Memory

### 27 Mar 2026 v1

@@ C.4 Properties of the Cubic Newton–Schulz Iteration
For $k=1$, the claim is clear. Assume Eq. (59) holds, then
$z(h^{(k+1)})^{\prime}(z)$ $=zh^{\prime}(h^{(k)}(z))(h^{(k)})^{\prime}(z)$
$=\frac{3}{2}\quantity(1-h^{(k)}(z)^{2})z(h^{(k)})^{\prime}(z)$
$\leq\frac{3}{2}\quantity(1-h^{(k)}(z)^{2})h^{(k)}(z).$
@@
Appendix D Proofs for Multiple Steps
D.1 Auxiliary Results
@@
Lemma 32 (Hanson-Wright inequality).
[…]
Proof.
See Vershynin [51, Theorem 6.2.2]. ∎
Lemma 33.
@@
Let ${\mathbf{Q}}:=\operatorname{diag}(q_{1},\dots,q_{N})$ and ${\mathbf{Z}}=\sqrt{d}\begin{bmatrix}u_{1}\;\cdots\;u_{N}\end{bmatrix}\in\mathbb{R}^{d\times N}$ so that $\mathbf{Z}$} has i.i.d. standard Gaussian entries and ${\mathbf{M}}=\frac{1}{d}{\mathbf{Z}}{\mathbf{Q}}{\mathbf{Z}}^{\top}$. By Lemma 30, it follows w.h.p. that
@@
Lemma 34 (superconcentration of Gaussian maxima).
@@
$\Pr(\max_{i\in[n]}Z_{i}\leq\sqrt{2\log 2n}-\frac{C\log\log n}{\sqrt{\log n}})=n^{-\omega(1)}.$
@@
$a_{n}=\sqrt{2\log 2n}-\frac{C\log\log n}{\sqrt{2\log 2n}}$
so that $a_{n}^{2}\leq 2\log 2n-2C\log\log n+1$ for sufficiently large $n$. It follows that
@@
$\leq\quantity(1-\frac{1}{2\sqrt{2\pi}a_{n}}\exp(-\log 2n+C\log\log n-\frac{1}{2}))^{n}$
$\leq\quantity(1-\frac{e^{C\log\log n}}{8n\sqrt{\pi e\log 2n}})^{n}$
$\leq\exp(-\frac{(\log n)^{C}}{8\sqrt{\pi e\log 2n}})$
@@ D.2 Proof of Theorem 9
We choose $\lambda_{0}=\lambda$ as in Theorem 3 and
@@
so that $\gamma_{t,ii}=\phi_{t}(q_{i}^{(t)})$. Repeating the argument in Lemma 12, we may express $\phi_{t}^{\prime}(0)$ in terms of the SVD of $\bar{{\mathbf{G}}}_{t,-i}$ via the Daleckii–Krein formula,
@@
Since $h_{t}(z)/z\leq\lambda_{t}^{-1}$ for all $z>0$, we immediately obtain the upper bound $\phi_{t}^{\prime}(0)\lesssim\lambda_{t}^{-1}$. For the lower bound, as in Lemma 17, we control the singular values of $\bar{{\mathbf{G}}}_{t,-i}$ using the eigenvalues of the corresponding weighted covariance matrix:
@@
When $t=0$, the bulk eigenvalue satisfies $\lambda_{d/2}({\mathbf{M}}_{0})\lesssim d^{-2\alpha}(\log d)^{2}\lesssim\lambda_{0}^{2}$ by Lemma 14. When $t\geq 1$, we instead use the following uniform bound.
Lemma 35.
@@
As in Lemma 14, we have that
@@
Moreover $\phi_{t}^{\prime}(q)\geq 0$ and $|\phi_{t}^{\prime\prime}(q)|\lesssim\lambda_{t}^{-2}$ hold identically as in Lemma 12 and Lemma 13. Expanding $\phi_{t}$ around zero, we obtain for all $t\leq\tau-1$,
@@
It remains to bound the interaction logits. For the first gradient step, we have shown in Proposition 18 that $u_{j}^{\top}h_{0}(\bar{{\mathbf{G}}}_{0})v_{i}$ is $\widetilde{O}(1/\sqrt{d})$ w.h.p. The interaction terms after the first step are much simpler to control; the Lipschitz concentration argument in Section B.8 will suffice. We only consider the case $j<i$ by symmetry. Fix $\tau$ such that $d_{\tau-1}<j\leq d_{\tau}$. By the same argument as Lemma 27, the map
@@ D.3 Proof of Theorem 10
Since $\sqrt{d}{\mathbf{U}}$ has i.i.d. standard Gaussian entries, by Lemma 30,
@@
Applying Lemma 30 again to $\sqrt{d}{\mathbf{V}}$ gives
@@
Since $B\gtrsim d_{\tau-1}^{\alpha}\log d\geq d_{t}^{\alpha}\log d$, we have that $\norm{q_{>d_{t}}}_{2}\lesssim d_{t}^{\frac{1}{2}-\alpha}\log d$ and $\lVert q_{>d_{t}}\rVert_{\infty}\leq d_{t}^{-\alpha}$ w.h.p. by Lemma 15. Plugging in above gives the desired result. ∎
Combining Lemma 31 and Lemma 36, we have that
@@
The argument for this regime is a slightly more involved version of Theorem 5. Consider a fixed item $i\asymp d^{\frac{1}{2\alpha}}d_{\tau}$ so that $i\lesssim B^{\frac{1}{\alpha}}(\log d)^{-\frac{1}{\alpha}}$ and competitors $j$ with $d_{\tau}<j\leq 2d_{\tau}$. Then $q_{j}^{(t)}/q_{i}^{(t)}\asymp p_{j}/p_{i}\asymp\sqrt{d}$ for all $t\leq\tau$ and
@@
If item $i$ was recovered at step $\tau+1$, it follows that $Z_{i}\geq Z_{j}-O(T)$ for all $j\in\mathcal{J}$. On the other hand, by Gaussian superconcentration (Lemma 34) it holds that $\max_{j\in\mathcal{J}}Z_{j}=\sqrt{2\log 2|\mathcal{J}|}+o(1)$. By Mill’s inequality, supposing $T=o(\sqrt{\log d})$,
$\Pr(Z_{i}\geq\sqrt{2\log 2|\mathcal{J}|}-O(T))\lesssim\frac{e^{-\frac{1}{2}\quantity(\sqrt{2\log 2|\mathcal{J}|}-O(T))^{2}}}{\sqrt{2\log 2|\mathcal{J}|}}\lesssim\frac{e^{O\quantity(T\sqrt{\log|\mathcal{J}|})}}{|\mathcal{J}|\sqrt{\log|\mathcal{J}|}}=\frac{o(\operatorname{poly}(d))}{\sqrt{d}}.$

### 28 Apr 2026 v2

@@ D.3 Properties of the Cubic Newton–Schulz Iteration
For $k=1$, the claim is clear. Assume Eq. (64) holds for $k\in\mathbb{N}$, then
$z(h^{(k+1)})^{\prime}(z)$ $=zh^{\prime}(h^{(k)}(z))(h^{(k)})^{\prime}(z)=\frac{3}{2}\quantity(1-h^{(k)}(z)^{2})z(h^{(k)})^{\prime}(z)\leq\frac{3}{2}\quantity(1-h^{(k)}(z)^{2})h^{(k)}(z).$
@@
Appendix E Proofs for Multiple Steps
E.1 Auxiliary Results
@@
Lemma E.3.
@@
Let ${\mathbf{Q}}:=\operatorname{diag}(q_{1},\dots,q_{N})$ and ${\mathbf{Z}}=\sqrt{d}\begin{bmatrix}u_{1}\;\cdots\;u_{N}\end{bmatrix}\in\mathbb{R}^{d\times N}$ so that $\mathbf{Z}$} has i.i.d. standard Gaussian entries and ${\mathbf{M}}=\frac{1}{d}{\mathbf{Z}}{\mathbf{Q}}{\mathbf{Z}}^{\top}$. By Lemma E.1, it follows w.h.p. that
@@
Lemma E.4 (superconcentration of Gaussian maxima).
@@
$\Pr(\max_{i\in[n]}Z_{i}\leq\sqrt{2\log n}-\frac{C\log\log n}{\sqrt{\log n}})=n^{-\omega(1)}.$
@@
$a_{n}=\sqrt{2\log n}-\frac{C\log\log n}{\sqrt{2\log n}}$
so that $a_{n}^{2}\leq 2\log n-2C\log\log n+1$ for sufficiently large $n$. It follows that
@@
$\leq\quantity(1-\frac{1}{2\sqrt{2\pi}a_{n}}\exp(-\log n+C\log\log n-\frac{1}{2}))^{n}$
$\leq\quantity(1-\frac{e^{C\log\log n}}{4n\sqrt{\pi e\log n}})^{n}$
$\leq\exp(-\frac{(\log n)^{C-0.5}}{4\sqrt{\pi e}})$
@@
We choose $\lambda_{0}=\lambda$ as in Theorem 4.1 and
@@
so that $\gamma_{t,ii}=\phi_{t}(q_{i}^{(t)})$. Repeating the argument in Lemma A.2, we may express $\phi_{t}^{\prime}(0)$ in terms of the SVD of $\bar{{\mathbf{G}}}_{t,-i}$ via the Daleckii–Krein formula,
@@
Since $h_{t}(z)/z\leq\lambda_{t}^{-1}$ for all $z>0$, we immediately obtain the upper bound $\phi_{t}^{\prime}(0)\lesssim\lambda_{t}^{-1}$. For the lower bound, as in Lemma A.7, we control the singular values of $\bar{{\mathbf{G}}}_{t,-i}$ using the eigenvalues of the corresponding weighted covariance matrix:
@@
When $t=0$, the bulk eigenvalue satisfies $\lambda_{d/2}({\mathbf{M}}_{0})\lesssim d^{-2\alpha}(\log d)^{2}\lesssim\lambda_{0}^{2}$ by Lemma A.4. When $t\geq 1$, we instead use the following uniform bound.
Lemma E.5.
@@
As in Lemma A.4, we have that
@@
Moreover $\phi_{t}^{\prime}(q)\geq 0$ and $|\phi_{t}^{\prime\prime}(q)|\lesssim\lambda_{t}^{-2}$ hold identically as in Lemma A.2 and Lemma A.3. Expanding $\phi_{t}$ around zero, we obtain for all $t\leq\tau-1$,
@@
It remains to bound the interaction logits. For the first gradient step, we have shown in Proposition B.1 that $u_{j}^{\top}h_{0}(\bar{{\mathbf{G}}}_{0})v_{i}$ is $\widetilde{O}(1/\sqrt{d})$ w.h.p. The interaction terms after the first step are much simpler to control; the Lipschitz concentration argument in Section B.8 will suffice. We only consider the case $j<i$ by symmetry. Fix $\tau$ such that $d_{\tau-1}<j\leq d_{\tau}$. By the same argument as Lemma B.10, the map
@@
Since $\sqrt{d}{\mathbf{U}}$ has i.i.d. standard Gaussian entries, by Lemma E.1,
@@
Applying Lemma E.1 again to $\sqrt{d}{\mathbf{V}}$ gives
@@
Since $B\gtrsim d_{\tau-1}^{\alpha}\log d\geq d_{t}^{\alpha}\log d$, we have that $\norm{q_{>d_{t}}}_{2}\lesssim d_{t}^{\frac{1}{2}-\alpha}\log d$ and $\lVert q_{>d_{t}}\rVert_{\infty}\leq d_{t}^{-\alpha}$ w.h.p. by Lemma A.5. Plugging in above gives the desired result. ∎
Combining Lemma E.2 and Lemma E.6, we have that
@@
The argument for this regime is a slightly more involved version of Theorem 4.3. Consider a fixed item $i\asymp d^{\frac{1}{2\alpha}}d_{\tau}$ so that $i\lesssim B^{\frac{1}{\alpha}}(\log d)^{-\frac{1}{\alpha}}$ and competitors $j$ with $d_{\tau}<j\leq 2d_{\tau}$. Then $q_{j}^{(t)}/q_{i}^{(t)}\asymp p_{j}/p_{i}\asymp\sqrt{d}$ for all $t\leq\tau$ and
@@
If item $i$ was recovered at step $\tau+1$, it follows that $Z_{i}\geq Z_{j}-O(T)$ for all $j\in\mathcal{J}$. On the other hand, by Gaussian superconcentration (Lemma E.4) it holds that $\max_{j\in\mathcal{J}}Z_{j}=\sqrt{2\log|\mathcal{J}|}+o(1)$. By Mill’s inequality, supposing $T=o(\sqrt{\log d})$,
$\Pr(Z_{i}\geq\sqrt{2\log|\mathcal{J}|}-O(T))\lesssim\frac{e^{-\frac{1}{2}\quantity(\sqrt{2\log|\mathcal{J}|}-O(T))^{2}}}{\sqrt{2\log|\mathcal{J}|}}\lesssim\frac{e^{O\quantity(T\sqrt{\log|\mathcal{J}|})}}{|\mathcal{J}|\sqrt{\log|\mathcal{J}|}}=\frac{o(\operatorname{poly}(d))}{\sqrt{d}}.$